OpenStudy (anonymous):

A rectangular pyramid has a base length, l, of 8.7cm, a base width, w, of 8.7cm, and a height, H, of 12.3cm. Find the volume of the pyramid to the nearest square unit. A. 306cm^3 B. 310cm^3 C. 315cm^3 D. 328cm^3

OpenStudy (anonymous):

Volume of pyramid = 1/3 * base area * pyramid's height = 1/3 * 8.7 * 8.7 *12.3 = 8.7 * 8.7 * 4.1 = 310.329 = 310 cm^3